OpenStudy (mtalhahassan2):

Mario borrows $4800 for 8.5 years at a fixed rate of simple interest. At the end of that time, he owes $8000. What interest rate is he being charged?

OpenStudy (jordanloveangel):

nterest = principal(rate)(time) He paid $8000-4800 - $3200 in interest 3200 = 4800(rate)(8.5) 3200 = 40,800r 0.078431372 or about 7.84%

OpenStudy (jordanloveangel):

Base Capital borrowed(C) = $4800 Time period for which the loan exists(n): 8.5 years Let the Rate of Interest(%) be= r (which is to be calculated) The Final amount(F=Base amount with interest) he woes after 8.5 years is $8000 So Total Interest (I) he paid= Final amount(F)-Base amount(C)= $8000-$4800= $3200 As per Formula of simple Interest total Interest(I) =C*r*n/100 => I = $4800* r * 8.5 /100= $3200 => r = ($3200*100) / ($4800*8.5) = 7.84% So rate of Interest = 7.84% per annum.

OpenStudy (jordanloveangel):

I = Prt r = I/Pr = 3200/(4800)(8.5) = 7.8%

OpenStudy (jordanloveangel):

hope this helps u

OpenStudy (mtalhahassan2):

ok thanks a lot

OpenStudy (jordanloveangel):

yeah np

OpenStudy (mtalhahassan2):

I have a question like for the formula I= Prt by u divide

OpenStudy (mtalhahassan2):

I = Prt r = I/Pr = 3200/(4800)(8.5) = 7.8%

OpenStudy (jordanloveangel):

https://answers.yahoo.com/question/index?qid=20090531123720AAn1MrJ look here this might help u lil